OpenStudy (anonymous):

3x + 10 4 3x + 10 4 @Mathematics

OpenStudy (anonymous):

\[\ge\]

hero (hero):

\[3x + 10 \ge 4\] subract 10 from both sides, then divide both sides by 3: \[x = -2\]

OpenStudy (karatechopper):

i think hero means \[x \ge -2\]